OpenStudy (anonymous):

1 2 3 4 5 6 7 8 9 | what is the minor of a22??

OpenStudy (kinggeorge):

If I'm correct about what a minor is, the minor of \(A_{22}\) would be the determinant of the matrix if you removed the second row and second column. In other words, the minor of \(A_{22}\) would then be \[(1*9)-(3*7) = 9-21=-12\]

OpenStudy (anonymous):

are minor and cofactor same thing??

OpenStudy (kinggeorge):

Close, but not quite. To get the cofactor, you would then have to multiply your minor by \((-1)^{i+j}\) where \(i, j\) are the indices of the row and column you're removing. In your case, they're both 2, so the cofactor would be \[(-1)^{2+2} * (-12) = (-1)^4 *(-12) = 1*(-12)=-12\]which is indeed the minor, but this is not always the case.